# A Cycle Ride to HDR: Semantics Aware Self-Supervised Framework for Unpaired LDR-to-HDR Image Translation

### v1

@@ 1 Introduction
High Dynamic Range (HDR) [4, 57] images capture a broader range of intensity values compared to Low Dynamic Range (LDR) or Standard Dynamic Range (SDR) counterparts, which have a pixel bit depth of only $2^{8}$ intensity levels. From human vision perspective, LDR images are often not visually pleasing, while from computer vision applications perspective, they hold limited information.
@@
Most commercial devices produce and display LDR content and some specialized hardware can support HDR, e.g., HDR cameras and sensors [37], which can capture images with more than $256$ intensity levels and HDR displays [12], which can display HDR intensity levels. Given that specialized hardware that captures HDR is expensive, researchers have been investigating approaches for accurate HDR reconstruction from LDR (i.e., inverse tone-mapping [78]) using conventional non-learning methods and modern data-driven approaches. For display purposes, however, one often needs to tone-map [27] the HDR content to LDR to fit the intensity range supported by standard hardware.
@@
Early attempts on HDR reconstruction address tasks such as image enhancement, e.g., recovering missing information due to extreme lighting conditions [62] or low lighting conditions [84, 47] and compression [11]. Current applications of HDR reconstruction span multiple domains, including, robotics/machine vision [83], media and entertainment [29], gaming, mixed reality and novel view synthesis [72, 53, 75, 31], as well as medical imaging [31].
@@
Early data-driven methods utilize Convolutional Neural Networks (CNN) [74], Transformers [73, 52] and Generative Adversarial Networks (GAN) [67, 63]. More recent approaches are based on Diffusion Models [15, 23], Neural Radiance Fields (NeRF) [31], and Gaussian Splatting [75]. Some methods use multi-exposed LDR [6, 71, 18, 48, 51, 45, 24, 90, 25, 69], while others use single-exposed LDR [5, 9, 88] paired with an HDR for supervised training.
@@
Although these methods achieve excellent results in reconstructing HDR images, most of them require proper {LDR,HDR} paired datasets [7, 51, 19] for training. Consequently, the quality of the state-of-the-art data-driven HDR reconstruction methods depends on the quality of the available paired {LDR,HDR} datasets. There is a research gap in the field; on the one hand, the literature on unpaired HDR reconstruction is extremely limited [49], and on the other, only a few approaches utilize semantic and contextual information to guide the reconstruction process [23, 22, 52, 77].
@@
To address this gap, we propose a method that leverages cycle consistency [89, 14] objective using unpaired data, where the model learns a mapping between domains, i.e., LDR $\leftrightarrow$ HDR. In addition, the method ensures semantic consistency between the LDR and reconstructed HDR by utilizing CLIP [66] embeddings and loss based on semantic segmentation. The proposed method reconstructs artifact-free and visually impressive HDR from single-exposed LDR images. It also outperforms the most recent state-of-the-art which predominantly uses paired datasets for training. Our work makes the following contributions:
@@
  • We propose novel generators based on modified U-Net architecture [70] that incorporates ConvLSTM-based feedback mechanism [30, 50] to mitigate visual artifacts in the HDR reconstruction (cf., Section 3.1).
  • We propose a CLIP embedding encoder for contrastive learning to minimize the semantic difference between LDR and reconstructed HDR pairs, while maximizing the difference between LDR and other {LDR,HDR} pairs (cf., Section 3.1).
  • We propose a novel loss function based on the Mean Intersection of Union (mIOU) metric to further ensure semantic consistency between the LDR and reconstructed HDR (cf., Section 3.2).
  • We perform thorough experimental validation for the contribution of all proposed components, both qualitatively and quantitatively (cf., Section 5).
@@ 2 Related Work
2.1 Non-Learning Approaches
The most popular approach in this category consists of multi-exposed LDR fusion to achieve high dynamic range in the output image [35]. This approach involves image feature alignment, calculating weights for feature mapping on the basis of image characteristics, and fusing the images on the basis of the weights to get the most appropriate exposures for each part of the image. Another approach is using histogram equalization [16] that allows for contrast and brightness levels re-adjustment in the over/underexposed regions of the image. Gradient domain manipulation [44] techniques enhance the granular details of an LDR image and expand its dynamic range. Pyramid-based image fusion [38] is a technique where image features are extracted into Laplacian pyramids [38] and then fused for the HDR. Some methods use Retinex-based [40, 58] tuning to effectively produce and display HDR images on consumer screens [40, 58]. Intensity Mapping Function based approaches [85, 17, 65] use transformations on the image pixels either by using Logarithmic mapping or Gamma correction to stretch the intensity range or enhance lower intensity pixels. The methods in this category, although time-efficient, face some issues including ghosting effects, artifacts, and blurring in the reconstruction, and they do not generalize well for variety of input LDR images.
2.2 Learning-Based Approaches
@@
The methods in this category either use single-exposed LDR as input [71, 18, 48, 51, 45, 24, 90, 25] or alternatively, multi-exposed LDR [63, 9, 69, 88, 6]. Barua et al. [5] harnessed the concept of histogram equalization of input LDR images, in addition to the original LDR, to overcome the contrast and hue miss-representation in over/underexposed areas of the LDR image. Cao et al. [10] proposed a channel-decoupled kernel-based approach which combines the output HDR with the output of another architecture in a pixel-wise fashion. Luzardo et al. [54] proposed a method to enhance the artistic intent of the reconstructed HDR by enhancing the peak brightness in the reconstruction process. In addition, recently, Neural Radiance Fields (NeRF) [60, 31, 59, 53], Diffusion Models [15, 23], and Gaussian Splatting [75] have also been used to improve the reconstruction, both in terms of space and time complexity.
@@
Some methods use weakly-supervised [45], self-supervised [88] and unsupervised [62] approaches for HDR reconstruction. Le et al. [45] proposed an indirect approach for HDR reconstruction from multi-exposed LDR using weak supervision. The method first outputs a stack of multi-exposed LDR, which are then merged using the state-of-the-art tool Photomatix [28] to obtain an HDR image. Zhang et al. [88] proposed a self-supervised approach for HDR reconstruction. The method requires multi-exposed LDR to learn the reconstruction network during training without the need for HDR counterparts. Nguyen et al. [62] proposed an unsupervised approach to recover image information from overexposed areas of LDR images. This approach does not require any HDR for supervision and instead uses pseudo ground truth images. Lee et al. [46] proposed a method with limited supervision for multi-exposed LDR generation consisting of a pair of networks that generate LDR with higher and lower exposure levels. Then the two exposure levels are combined to generate the HDR image.
@@
GAN-Based Approaches. Generative Adversarial Networks (GAN) [21, 3] architectures are well-explored in this domain. Niu et al. [63] proposed a GAN-based method that can handle images with foreground motions by fusing multi-exposed LDR and extracting important information from the over/underexposed areas of the LDR. Raipurkar et al. [67] proposed a conditional GAN architecture that adds details to the saturated regions of the input LDR using a pre-trained segmentation model to extract exposure masks. Guo et al. [24] proposed a two-stage pipeline that extracts the over/underexposed features with high accuracy. GAN with attention mechanism first generates the missing information in those extreme exposure areas, and in the second stage, a CNN with multiple branches fuses the multi-exposed LDR from the previous stage to reconstruct the HDR. Nam et al. [61] proposed a GAN method that uses exposure values for conditional generation of multi-exposure stack that adapts well to varying color and brightness levels.
@@
Li et al. [49] proposed a GAN-based method for unpaired multi-exposure LDR to HDR translation. The method introduces modified GAN loss and novel discriminator to tackle ghosting artifacts caused from the misalignment in the LDR stack and HDR. Wu et al. [82] proposed CycleGAN-like architecture for low-light image enhancement task using unpaired data for training. The method fuses the Retinex theory with CycleGAN concept to enhance the lighting conditions globally, recover color and reduce noise in the output HDR image.
@@
Table 1: Summary of recent state-of-the-art methods. I/O: LDR used as input (SE: Single-exposed and ME: Multi-exposed), O/P: Reconstructs directly HDR (D) or multi-exposed LDR stack (I), Unpaired: Uses unpaired data, Context: Uses local/global image information and relationship among entities in the image, Semantics: Uses color/texture information and identity of the items in the image, Artifacts: Handles visual artifacts, TM: Also performs tone-mapping i.e. HDR $\rightarrow$ LDR.
Method I/O O/P Unpaired Context Semantics Artifacts TM
@@
Ours SE D
Semantic and Knowledge-Based Approaches. Some methods attempt to use semantic/context information in the image or image formation process for HDR reconstruction. Wang et al. [77] proposed a method which approximates the inverse of the camera pipeline. Their knowledge-inspired block uses the image formation knowledge to address three tasks during HDR reconstruction: missing details recovery, image parameters adjustment, and image noise reduction. Goswami et al. [23] proposed a method to recover the clipped intensity values of an LDR/SDR image due to the tone-mapping process in the camera. The proposed method works in two stages - first it uses a semantic graph-based guidance to help the diffusion process with the in-painting of saturated image parts, and second, the problem is formulated as HDR in-painting from SDR in-painted regions. Liu et al. [52] proposed a vision transformer approach with context-awareness to remove ghosting effects in the output HDR. The model is a dual-branch architecture to capture both local and global context in the input image that enables the generation process to remove unwanted information in the output HDR and avoid artifacts.
Limitations. Our analysis reveals that methods based on single-exposed LDR fail to preserve image characteristics such as color hue and saturation, color contrast, and intensity brightness levels in the HDR. On the other hand, methods based on multi-exposed LDR produce unwanted visual artifacts, missing details and ghosting effects. While most methods produce excellent results in terms of visual quality, the main limitations of these approaches are the requirement of paired {LDR,HDR} datasets for training, and the lack of semantic and contextual knowledge guidance in the reconstruction process which could significantly improve the quality of output HDR images. Table 1 summarizes a comparison between recent state-of-the-art and our method in terms of various parameters.
Figure 1: Overview of the proposed method architecture (cf., Section 3.1) where $x$ and $y$ represent LDR and HDR images, respectively. The method is trained with five objectives: adversarial, cycle consistency, identity, contrastive, and semantic segmentation (cf., Section 3.2).
@@ 3.1 Architecture Modules
We adopt a CycleGAN [89] architecture as the basis of our method depicted in Figure 1. The network includes two generators and two discriminators (cf., Supplementary for brief background on cycle consistency). Let us denote the two domains as $X$ for LDR images and $Y$ for HDR images. Furthermore, let us denote the generator used in the forward cycle that maps images from LDR to HDR as $G_{Y}$ and the one used in the backward cycle that maps images from HDR to LDR as $G_{X}$. Finally, let us denote the discriminator that discriminates between the reconstructed LDR and real LDR images as $D_{X}$ and the one that discriminates between the reconstructed HDR and real HDR images as $D_{Y}$. The images from the two domains can be represented as $\{x_{i}\}_{i=1}^{N}$ where $x_{i}\in X$ and $\{y_{j}\}_{j=1}^{M}$ where $y_{j}\in Y$. The data distribution of the two domains can be represented as $x\sim p_{\text{data}}(x)$ and $y\sim p_{\text{data}}(y)$.
@@
Generators. The generators $G_{Y}$ and $G_{X}$ are based on U-Net architecture [70] that includes an encoder and a decoder block with skip connections from each level of the encoder to the decoder. In our generators, we propose a feedback mechanism [30, 86] between the encoder and decoder block. The rationale behind the feedback is to refine the features extracted from the encoder (during the first iteration of the feedback) to guide the decoder for better output image reconstruction over the rest of the iterations. Hence, the feedback block not only iterates over its own output but also re-runs the decoder situated ahead of it in each iteration while keeping the encoder frozen until the feedback iteration completes. The feedback is implemented with a ConvLSTM [30, 50] network and the number of iterations is fixed to $4$ based on an ablation experiment testing $2$, $3$, and $4$ iterations setups (cf., Supplementary for visual depiction of the proposed generators and implementation details).
Discriminators. The discriminators $D_{X}$ and $D_{Y}$ are based on [49, 32] (cf., Supplementary for implementation details).
@@
Encoders. We also introduce a CLIP embeddings encoder $E$. Specifically we use the pre-trained Vision Transformer (ViT) [2] CLIP encoder from [66] which helps in extracting image embeddings with both local and global semantic context. For the forward cycle consistency we use $E(x)$ and $E(y)$. Then we add the embeddings from the encoder and feed them back to the bottleneck layer of the $G_{Y}$ decoder. Similarly, for the backward cycle consistency we add the embeddings from $E(y)$ and $E(x)$ and feed them back to the bottleneck layer of the $G_{X}$ decoder.

### v2

# LLM-HDR: Bridging LLM-based Perception and Self-Supervision for Unpaired LDR-to-HDR Image Reconstruction

@@ Abstract
Figure 1: Qualitative comparison of the proposed LLM-HDR method and state-of-the-art method SelfHDR [116] (ICLR’24). We see, our method handles the overexposed portions in the sky more realistically.
@@ 1 Introduction
High Dynamic Range (HDR) [3, 70] images capture a wider range of intensity values compared to their Low Dynamic Range (LDR) or Standard Dynamic Range (SDR) counterparts, which have a pixel bit depth of only $2^{8}$ intensity levels. From human vision perspective, LDR images are often not visually pleasing, while from the computer/robot vision applications perspective, they hold limited information.
@@
Most commercial devices produce and display LDR content, and some specialized hardware can support HDR, e.g., HDR cameras and sensors [45], which can capture images with more than $256$ intensity levels, and HDR displays [14], which can display HDR intensity levels. Given that specialized hardware that captures HDR is expensive, researchers have been investigating approaches for accurate HDR reconstruction from LDR (i.e., inverse tone-mapping [98]) using conventional non-learning methods and modern data-driven approaches. For display purposes, however, one often needs to tone-map [33] the HDR content to LDR to fit the intensity range supported by standard hardware.
@@
Early HDR reconstruction methods address tasks such as image enhancement, e.g., recovering missing information due to extreme lighting conditions [76] or low lighting conditions [109, 57] and compression [13]. Current applications of HDR reconstruction span multiple domains, including, robotics/machine vision [103], media and entertainment [35], gaming, mixed reality, and novel view synthesis [88, 66, 91, 37], as well as medical imaging [37].
@@
Early data-driven methods utilize Convolutional Neural Networks (CNN) [90], Transformers [89, 65] and Generative Adversarial Networks (GAN) [82, 77]. Recent approaches for HDR image and/or 3D HDR scene reconstruction are based on Diffusion Models [19, 28, 8, 95], Neural Radiance Fields (NeRF) [37], and Gaussian Splatting [91]. Some use multi-exposed [5, 87, 22, 60, 64, 55, 29, 118, 30, 85], while others use single-exposed LDR [6, 11, 116] paired with HDR images for supervised training.
Large Language Models (LLM) [75] and Vision-Language Models (VLM) [114, 105] have been widely used in many vision tasks, e.g., image generation [92], deepfake detection [41], semantic segmentation [97], image relighting [15], and adverse weather condition removal [108]. However, the human-defined knowledge of such models has not been utilized for LDR-to-HDR image reconstruction.
@@
Although current methods achieve excellent results in reconstructing HDR images, most of them require proper {LDR,HDR} paired datasets [7, 64, 23] for training. Consequently, the quality of the state-of-the-art data-driven HDR reconstruction methods depends on the quality of the available paired {LDR,HDR} datasets. There is a research gap in the field; on the one hand, the literature on unpaired HDR reconstruction is extremely limited [61, 94], and on the other, only a few approaches utilize semantic and contextual information to guide the reconstruction [28, 27, 65, 96].
@@
To address this gap, we propose LLM-HDR, a method that leverages LLM-driven cycle consistency [117, 17] objective using unpaired data, where the model learns a mapping between domains, i.e., LDR $\leftrightarrow$ HDR. In addition, the method ensures semantic consistency between the LDR and reconstructed HDR by utilizing Contrastive Language-Image Pretraining (CLIP) [81] embeddings and loss based on semantic segmentation. Moreover, LLM-based artifact- and exposure-aware information further supports the training process to deliver more realistic and natural HDR image generation. The proposed method reconstructs artifact-free and visually impressive HDR from single-exposed LDR images. It also outperforms the most recent state-of-the-art which predominantly uses paired datasets for training. Our work makes the following contributions:
@@
  • We leverage LLM-based loss and artifact- and exposure-aware saliency maps using human-defined prompting to further refine those areas in the reconstructed HDR images (Secs. 3.1 and 3.2).
  • We propose a novel generator based on a modified U-Net architecture [86] that incorporates ConvLSTM-based artifact-aware feedback mechanism [36, 63] and exposure-aware skip connections to mitigate visual artifacts in the HDR reconstruction (Sec. 3.1).
  • We propose a CLIP embedding encoder for contrastive learning to minimize the semantic difference between LDR and reconstructed HDR image pairs, while maximizing the difference between LDR and other {LDR,HDR} image pairs (Sec. 3.1).
  • We propose a novel loss function based on the Mean Intersection over Union (mIoU) metric to further ensure semantic consistency between the LDR and reconstructed HDR images (Sec. 3.2).
  • We perform thorough experimental validation for the contribution of all proposed components, both qualitatively and quantitatively (Sec. 5).
@@ 2 Related Work
Learning-based approaches use either single-exposed LDR as input [87, 22, 60, 64, 55, 29, 118, 30] or alternatively, multi-exposed LDR [77, 11, 85, 116, 5]. Some of the multi-exposed methods use novel feature fusion or aggregation methods [111, 104, 107]. Barua et al. [6] harnessed the concept of histogram equalization of input LDR, in addition to the original LDR, to overcome the contrast and hue miss-representation in over/underexposed areas of the LDR. Cao et al. [12] proposed a channel-decoupled kernel-based approach which combines the output HDR with the output of another architecture in a pixel-wise fashion. Luzardo et al. [67] proposed a method to enhance the artistic intent of the reconstructed HDR by enhancing the peak brightness in the reconstruction process. In addition, recently, Neural Radiance Fields [72, 37, 73, 66], Diffusion Models [19, 28, 95], and Gaussian Splatting [91] have also been used to improve the reconstruction of HDR content.
@@
Some methods use weakly-supervised [55], self-supervised [116] and unsupervised [76, 94] approaches for HDR reconstruction. Le et al. [55] proposed an indirect approach for HDR reconstruction from multi-exposed LDR using weak supervision. The method first outputs a stack of multi-exposed LDR, which are then merged using the state-of-the-art tool Photomatix [34] to obtain an HDR. Zhang et al. [116] proposed a self-supervised approach for HDR reconstruction. The method requires multi-exposed LDR to learn the reconstruction network during training without the need for HDR counterparts. Nguyen et al. [76] proposed an unsupervised approach to recover image information from overexposed areas of LDR. This approach does not require any HDR for supervision and instead uses pseudo-ground truth images. Lee et al. [56] proposed a method with limited supervision for multi-exposed LDR generation consisting of a pair of networks that generate LDR with higher and lower exposure levels. Then the two exposure levels are combined to generate the HDR. Wang et al. [94] proposed an unsupervised approach for HDR reconstruction. They train the model in such a way that when the HDR is re-projected to LDR using a camera response function model it becomes indistinguishable from the original LDR ( Sec. S1 for related works on non-learning methods).
@@
GAN-Based Approaches. Generative Adversarial Networks [26, 2, 94] are well-explored in this domain. Niu et al. [77] proposed a GAN-based method that can handle images with foreground motions by fusing multi-exposed LDR and extracting important information from the over/underexposed areas of the LDR. Raipurkar et al. [82] proposed a conditional GAN architecture that adds details to the saturated regions of the input LDR using a pre-trained segmentation model to extract exposure masks. Guo et al. [29] proposed a two-stage pipeline that extracts the over/underexposed features with high accuracy. GAN with attention mechanism first generates the missing information in those extreme exposure areas, and in the second stage, a CNN with multiple branches fuses the multi-exposed LDR from the previous stage to reconstruct the HDR. Nam et al. [74] proposed a GAN method that uses exposure values for conditional generation of multi-exposure stack that adapts well to varying color and brightness levels.
@@
Li et al. [61] proposed a GAN-based unsupervised method for unpaired multi-exposure LDR-to-HDR translation. The method introduces modified GAN loss and a novel discriminator to tackle ghosting artifacts caused from the misalignment in the LDR stack and HDR. Wu et al. [102] proposed CycleGAN-like architecture for low-light image enhancement tasks using unpaired data for training. The method fuses the Retinex theory [54] with CycleGAN [117] concept to enhance the lighting conditions globally, recover color and reduce noise in the output HDR.
Semantic and Knowledge-Based Approaches. Some methods attempt to use semantic/context information in the image or image formation process for HDR reconstruction. Wang et al. [96] proposed a method that approximates the inverse of the camera pipeline. Their knowledge-inspired block uses the knowledge of image formation to address three tasks during HDR reconstruction: missing details recovery, adjustment of image parameters, and reduction of image noise. Goswami et al. [28] proposed a method to recover the clipped intensity values of an LDR/SDR image due to the tone-mapping process in the camera. The proposed method works in two stages: first it uses a semantic graph-based guidance to help the diffusion process with the in-painting of saturated image parts, and second, the problem is formulated as HDR in-painting from SDR in-painted regions. Liu et al. [65] proposed a vision transformer approach with context-awareness to remove ghosting effects in the output HDR. The model is a dual-branch architecture to capture both local and global context in the input image that enables the generation process to remove unwanted information in the output HDR and avoid artifacts.
Limitations. Our analysis reveals that the methods based on single-exposed LDR fail to preserve image characteristics such as the levels of color hue and saturation, contrast, and brightness intensity in HDR. On the other hand, methods based on multi-exposed LDR produce unwanted visual artifacts such as halo in edges and boundaries, outlier pixel intensities, unwanted repeated pattens, missing details in shadow and highly bright regions, irregular color and texture transitions, and ghosting effects from dynamic scenes. While most methods produce excellent visual quality, the main limitations of these approaches are the requirement of paired {LDR,HDR} datasets for training, and the lack of semantic and contextual knowledge guidance in the reconstruction process that could significantly improve the quality of HDR output. Finally, LLM- and VLM-based [114, 105] perception capabilities have not been utilized in this problem area, which might improve performance. Tab. 1 summarizes a comparison between the state-of-the-art and our method in terms of various parameters.
@@
Table 1: Summary of recent state-of-the-art methods. I/O: LDR used as input, single-exposed (SE) and multi-exposed (ME), O/P: Reconstructs directly HDR (D) or multi-exposed LDR stack (I), UP: Can be trained with unpaired data, LLM: Uses LLM-based perception, Context: Uses local/global image information and relationship among entities in the image, Semantics: Uses color/texture information and identity of the items in the image, Artifacts: Handles visual artifacts in heavily over/underexposed areas, TM: Also performs tone-mapping i.e. HDR $\rightarrow$ LDR.
Method I/O O/P UP LLM Context Semantics Artifacts TM
@@
LLM-HDR (Ours) SE D
Figure 2: Overview of the proposed method architecture (Sec. 3.1, Sec. S2.2) where $x$ and $y$ represent LDR and HDR images, respectively. The method is trained with six objectives: adversarial, cycle consistency, identity, LLM-based, contrastive, and semantic segmentation (Sec. 3.2, Sec. S2.3).
@@ 3.1 Architecture Modules
We adopt an cycle-consistent adversarial [117] architecture as the basis of our method depicted in Fig. 2. The network includes two generators and two discriminators. The network also takes advantage of CLIP encoders and LLM-based perception to further bridge the gap between semantic and perceptual differences between input and generated images (Sec. S2.1 for a concise background on the key concepts employed in our method). Let us denote the two domains as $X$ for LDR images and $Y$ for HDR images. Furthermore, let us denote the generator used in the forward cycle that maps images from LDR to HDR as $G_{Y}$ and the one used in the backward cycle that maps images from HDR to LDR as $G_{X}$. Finally, let us denote the discriminator that discriminates between the reconstructed LDR and real LDR images as $D_{X}$ and the one that discriminates between the reconstructed HDR and real HDR images as $D_{Y}$. The images from the two domains can be represented as $\{x_{i}\}_{i=1}^{N}$ where $x_{i}\in X$ and $\{y_{j}\}_{j=1}^{M}$ where $y_{j}\in Y$. The data distribution of the two domains can be represented as $x\sim p_{\text{data}}(x)$ and $y\sim p_{\text{data}}(y)$.
@@
Generators. The generators $G_{Y}$ and $G_{X}$ are based on U-Net architecture [86] that includes an encoder and a decoder block with skip connections from each level of the encoder to the decoder. In our generators, we propose a feedback mechanism [36, 113] between the encoder and decoder block. The rationale behind the feedback is to refine the features extracted from the encoder (during the first iteration of the feedback) to guide the decoder for better output image reconstruction over the rest of the iterations. Hence, the feedback block not only iterates over its own output but also re-runs the decoder situated ahead of it in each iteration while keeping the encoder frozen until the feedback iteration completes. The feedback is implemented with a ConvLSTM [36, 63] network and the number of iterations is fixed to $4$ based on an ablation experiment testing $2$, $3$, and $4$ iterations setups (Sec. S2.2 for visual depiction of the proposed generators and implementation details).
Discriminators. The discriminators $D_{X}$ and $D_{Y}$ are based on [61, 39] (Sec. S2.2 for implementation details).
LLM. We adopt a pre-trained LLM [38] for zero-shot Q&A sessions (using human-defined knowledge) in each iteration of the training process. We prompt the LLM to provide the pixels identified as artifacts, and over/underexposed regions for both the reconstructed HDR and LDR. Given that these three factors constitute the nature of the LDR $\leftrightarrow$ HDR translation, the pixels proposed by the LLM are used in a LLM-based loss function. In addition, the LLM generates a separate saliency map for each of those areas, i.e., artifact, overexposed, and underexposed maps. The artifact map is multiplied with the input features to the feedback mechanism of the generators $G_{Y}$ and $G_{X}$. The underexposed map is fed into each of the skip connections (levels $1$ to $3$) in the generator $G_{Y}$ only and the overexposed map is multiplied with the input features to the bottleneck layer of this generator. We use element-wise multiplication as a gating mechanism to emphasize the selected areas in the maps. The rationale behind the separated saliency map fusion is based on the proposed U-Net architecture where distorted pixels are mostly refined by the feedback mechanism and later layers, high-intensity lighting details are generally recovered by the middle or later layers, and low-intensity shadows are reconstructed realistically at the starting and ending layers (Sec. S2.2 for visual depiction of the proposed LLM module and implementation details).
@@
Encoders. We also introduce a CLIP embeddings encoder $E$. Specifically, we use a pre-trained CLIP encoder [81] to extract image embeddings with both local and global semantic context. For the forward cycle consistency we use $E(x)$ and $E(\hat{y})$. Then we add the embeddings and feed them back to the bottleneck layer of the $G_{Y}$ decoder. Similarly, for the backward cycle consistency we add the embeddings from $E(y)$ and $E(\hat{x})$ and feed them back to the bottleneck layer of the $G_{X}$ decoder (Sec. S2.2 for implementation details).